# Learning with Incomplete Context: Linear Contextual Bandits with Pretrained Imputation

### v2

@@ Appendix G Setup of the Lower Bound
$\mathcal{K}^{(\texttt{non})}_{\bm{\theta},t}(f,f^{\prime})$ $=C_{0}\sum_{i=1}^{N}\mathbb{E}_{f}\left[f\left(\bm{O}_{i}^{(0)}\right)-f^{\prime}\left(\bm{O}_{i}^{(0)}\right)\right]^{2}$ (H.1)
$~~+C_{D}\sum_{\tau=1}^{t}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\left(f(\bm{O}_{\tau})-f^{\prime}(\bm{O}_{\tau})\right)^{2}\mathbbm{1}\left\{A_{\tau}=1,V_{\tau}=0\right\}\mid\mathcal{H}_{\tau-1}\right]$
@@
$\mathcal{K}^{(\texttt{lin})}_{f,t}\left(\bm{\theta},\bm{\theta}^{\prime}\right):=C_{D}\sum_{\tau=1}^{t}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathbbm{1}(V_{\tau}=1)\left(\bm{\theta}_{Q}^{\top}\bm{Q}_{\tau}^{(A_{\tau})}-\bm{\theta}^{{}^{\prime}\top}_{Q}\bm{Q}_{\tau}^{(A_{\tau})}\right)^{2}\mid\mathcal{H}_{\tau-1}\right].$ (H.2)
Then for any fixed policy $\pi$ and distribution $\mathbb{P}^{(t)}_{\bm{\theta},f,\pi}$ whose density is specified in Equation (G.16),
$\operatorname{KL}\left(\mathbb{P}^{(t)}_{\bm{\theta},f,\pi}\|\mathbb{P}^{(t)}_{\bm{\theta}^{\prime},f^{\prime},\pi}\right)\leq\mathcal{K}^{(\texttt{non})}_{\bm{\theta},t}(f,f^{\prime})+\mathcal{K}^{(\texttt{lin})}_{f,t}\left(\bm{\theta},\bm{\theta}^{\prime}\right).$ (H.3)
The proof of all the technical lemmas are deferred to Section H.1. Lemma H.2 controls $\mathcal{K}^{(\texttt{lin})}_{f,t}$, while $\mathcal{K}^{(\texttt{non})}_{\bm{\theta},t}(f,f^{\prime})$ is controlled by Equation (H.31) in the proof of Lemma H.5.
Lemma H.2.
@@
$\mathbb{P}\left(\bm{Q}=\mathbf{e}_{i}\right)=\frac{1}{d_{\texttt{lin}}}\quad\text{for }i\in[d_{\texttt{lin}}].$ (H.4)
@@
$\mathcal{K}^{(\texttt{lin})}_{f,t}\left(\bm{\theta},\bm{\theta}^{\prime}\right)=\frac{C_{D}t\left\|\bm{\theta}_{Q}-\bm{\theta}^{\prime}_{Q}\right\|_{2}^{2}}{2d_{\texttt{lin}}}.$ (H.5)
@@
For the distribution in Equation (G.16), the expected cumulative regret is given by
$\mathcal{R}_{T}(\bm{\theta},f)$ $=\sum_{t=1}^{T}\mathbb{E}^{(t-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[R^{\star}_{t}-R_{t}\mid\mathcal{H}_{t-1}\right]$ (H.6)
$=\sum_{t=1}^{T}\mathbb{E}^{(t-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathbbm{1}\left(V_{t}=1\right)\left(\bm{Q}^{\star}_{t}-\bm{Q}^{(A_{t})}_{t}\right)^{\top}\bm{\theta}_{Q}\mid\mathcal{H}_{t-1}\right]$
$~~+\frac{1}{2}\sum_{t=1}^{T}\mathbb{E}^{(t-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathbbm{1}\left(V_{t}=0\right)\left(f^{\star}(\bm{O}_{t})-f^{(A_{t})}(\bm{O}_{t})\right)\mid\mathcal{H}_{t-1}\right]$
@@
Lemma H.3 controls $\mathcal{R}_{T}^{(\texttt{lin})}(\bm{\theta},f)$.
Lemma H.3.
@@
$\sup_{\theta\in\Theta}\mathcal{R}^{(\texttt{lin})}_{T}(\bm{\theta},f)$ $\geq\frac{\sqrt{d_{\texttt{lin}}T}}{4}\exp\left\{-2C_{D}\right\}.$ (H.7)
@@
where $M>0$ is specified later in Equation (H.45). We index the bins by integers $k\in[M^{d_{\texttt{non}}}]$ via the mapping
@@
$\phi_{\beta}(\mathbf{o})=\begin{cases}(1-\|\mathbf{o}\|_{\infty})^{\beta}&\mbox{ if }0\leq\|\mathbf{o}\|_{\infty}\leq 1,\\ 0&\mbox{ if }\|\mathbf{o}\|_{\infty}>1.\end{cases}$ (H.8)
@@
$m=\lceil c_{m}M^{d_{\texttt{non}}}\rceil$ (H.9)
@@
$f_{\bm{\omega}}(\mathbf{o})=\sum_{j=1}^{m}\omega_{j}\varphi_{j}(\mathbf{o}),$ (H.10)
@@
$\varphi_{j}(\mathbf{o})=M^{-\beta}C_{\phi}\phi_{\beta}\left(2M[\mathbf{o}-\bm{b}_{j}]\right)\mathbbm{1}(\mathbf{o}\in B_{j})$ (H.11)
and $C_{\phi}>0$ is a constant specified in Equation (H.12). Note that for any $\mathbf{o}\in B_{j}$, the rescaled argument satisfies $2M(\mathbf{o}-\bm{b}_{j})\in\left[-1,1\right]^{d_{\texttt{non}}}$, so $\left\|2M(\mathbf{o}-\bm{b}_{j})\right\|_{\infty}\in[0,1]$, ensuring that $\varphi_{j}$ in Equation (H.11) is well-defined.
The function $f_{\bm{\omega}}$ is thus a linear combination of localized, smooth bump functions with disjoint supports. Lemma H.4 establishes that each $f_{\bm{\omega}}$ lies in $\mathcal{F}_{\beta,L}$ for a suitable choice of constant $L$.
Lemma H.4.
Suppose that $\beta\in(0,1]$. For any $\bm{\omega}\in\Omega_{m}=\{\pm 1\}^{m}$, the function $f_{\bm{\omega}}$ defined in Equation (H.10) belongs to the smoothness class $\mathcal{F}_{\beta,L}$ with $L=\beta 2^{\beta}C_{\phi}>0$.
@@
$C_{\phi}:=\frac{L}{\beta 2^{\beta}}.$ (H.12)
Based on this choice of packing set, Lemma H.5 controls $\mathcal{R}_{T}^{(\texttt{non})}(\bm{\theta},f)$.
Lemma H.5.
@@
$\sup_{f\in\mathcal{F}_{\beta,L}}\mathcal{R}^{(\texttt{non})}_{T}(\bm{\theta},f)=\Theta\left(TN^{-\frac{\beta}{2\beta+d_{\texttt{non}}}}\right),$ (H.13)
where $\mathcal{R}^{(\texttt{non})}$ is defined in Equation (H.6).
Taking Equations (H.13) with (H.19) into Equation (H.6), we have
@@
establishing the desired result in Equation (G.17).
H.1 Proof of Technical Lemmas
In this section, we present the proof of the Lemmas H.1-H.5 used in the proof of Theorem G.1. We will frequently use the Bretagnolle-Huber inequality given in the following theorem.
Theorem H.1 (Bretagnolle-Huber inequality).
@@ H.1 Proof of Technical Lemmas
See Theorem 14.2 in Lattimore and Szepesvári, (2020). ∎
H.1.1 Proof of Lemma H.1
@@
Recall the definition of $\mathbb{P}^{(t)}_{\bm{\theta},f,\pi}$ as stated in Equation (G.16). Eliminating the shared terms, it follows that
$\operatorname{KL}\left(\mathbb{P}^{(t)}_{\bm{\theta},f,\pi}\|\mathbb{P}^{(t)}_{\bm{\theta}^{\prime},f^{\prime},\pi}\right)$ $=\mathbb{E}^{(t)}_{\bm{\theta},f,\pi}\left[\log\frac{d\mathbb{P}^{(t)}_{\bm{\theta},f,\pi}}{d\mathbb{P}^{(t)}_{\bm{\theta}^{\prime},f^{\prime},\pi}}\right$] (H.14)
@@
For $\mathcal{K}_{1}$ in Equation (H.14), by the KL divergence assumption in Equation (G.13), we have
$\mathcal{K}_{1}=\sum_{i=1}^{N}\mathbb{E}_{f}\left[\log\frac{p^{(0)}_{f(\bm{O}^{(0)}_{i})}}{p^{(0)}_{f^{\prime}(\bm{O}^{(0)}_{i})}}\left(W^{(0)}_{i}\right)\right]\leq C_{0}\sum_{i=1}^{N}\mathbb{E}_{f}\left[f\left(\bm{O}^{(0)}_{i}\right)-f^{\prime}\left(\bm{O}^{(0)}_{i}\right)\right]^{2}.$ (H.15)
@@
$=\mathbb{E}^{(t)}_{\bm{\theta},f,\pi}\left[\log\frac{p_{\bm{\theta}^{\top}_{Q}\bm{Q}_{t}^{(A_{t})}}\left(R_{t}\right)}{p_{\bm{\theta}^{\prime\top}_{Q}\bm{Q}_{t}^{(A_{t})}}\left(R_{t}\right)}\mathbbm{1}\{V_{t}=0\}\right]+\mathbb{E}^{(t)}_{\bm{\theta},f,\pi}\left[\log\frac{p_{f^{(A_{t})}(\bm{O}_{t})}\left(R_{t}\right)}{p_{f^{\prime(A_{t})}(\bm{O}_{t})}\left(R_{t}\right)}\mathbbm{1}\{V_{t}=1\}\right]$
$=\sum_{a\in\mathcal{A}}\mathbb{E}^{(t-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathbbm{1}(A_{t}=a,V_{t}=0)\operatorname{KL}\left(\mathbb{P}_{\bm{\theta}_{Q}^{\top}\bm{Q}_{t}^{(a)}}\|\mathbb{P}_{\bm{\theta}^{{}^{\prime}\top}_{Q}\bm{Q}_{t}^{(a)}}\right)\mid\mathcal{H}_{t-1}\right]$
$~~+\sum_{a\in\mathcal{A}}\mathbb{E}^{(t-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathbbm{1}(A_{t}=a,V_{t}=1)\operatorname{KL}\left(\mathbb{P}_{f^{(a)}(\bm{O}_{t})}\|\mathbb{P}_{f^{\prime(a)}(\bm{O}_{t})}\right)\mid\mathcal{H}_{t-1}\right],$
where the last equality follows from the definition of KL divergence. Taking the above display and Equation (H.15) into Equation (H.14) yields that
$\operatorname{KL}\left(\mathbb{P}^{(t)}_{\bm{\theta},f,\pi},\mathbb{P}^{(t)}_{\bm{\theta}^{\prime},f^{\prime},\pi}\right)$ $=\mathcal{K}_{1}+\sum_{\tau=1}^{t}\sum_{a\in\mathcal{A}}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathbbm{1}(A_{\tau}=a,V_{\tau}=0)\operatorname{KL}\left(\mathbb{P}_{f^{(a)}(\bm{O}_{\tau})}\|\mathbb{P}_{f^{\prime(a)}(\bm{O}_{\tau})}\right)\mid\mathcal{H}_{\tau-1}\right]$ (H.16)
$+\sum_{\tau=1}^{t}\sum_{a\in\mathcal{A}}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathbbm{1}(A_{\tau}=a,V_{\tau}=1)\operatorname{KL}\left(\mathbb{P}_{\bm{\theta}_{Q}^{\top}\bm{Q}_{\tau}^{(a)}}\|\mathbb{P}_{\bm{\theta}^{{}^{\prime}\top}_{Q}\bm{Q}_{\tau}^{(a)}}\right)\mid\mathcal{H}_{\tau-1}\right]$
@@
$+C_{D}\sum_{\tau=1}^{t}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\left(f(\bm{O}_{\tau})-f^{\prime}(\bm{O}_{\tau})\right)^{2}\mathbbm{1}\left\{A_{\tau}=1,V_{\tau}=0\right\}\mid\mathcal{H}_{\tau-1}\right]$
$+C_{D}\sum_{\tau=1}^{t}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathbbm{1}(V_{\tau}=1)\left(\bm{\theta}_{Q}^{\top}\bm{Q}_{\tau}^{(A_{\tau})}-\bm{\theta}^{{}^{\prime}\top}_{Q}\bm{Q}_{\tau}^{(A_{\tau})}\right)^{2}\mid\mathcal{H}_{\tau-1}\right],$
where the last inequality follows from Assumption G.1 and Equation (G.15). Taking the definition of $\mathcal{K}^{(\texttt{non})}_{\bm{\theta},t}$ and $\mathcal{K}^{(\texttt{lin})}_{f,t}$ in Equations (H.1) and (H.2) into Equation (H.16) yields the desired bound in Equation (H.3). ∎
H.1.2 Proof of Lemma H.2
@@
By definition of $\mathbb{P}_{Q}$ in Equation (H.4),
@@
$\mathbb{E}\left[\mathbbm{1}(V_{t}=1)\left(\bm{\theta}_{Q}^{\top}\bm{Q}_{t}^{(A_{t})}-\bm{\theta}^{{}^{\prime}\top}_{Q}\bm{Q}_{t}^{(A_{t})}\right)^{2}\mid\mathcal{H}_{t-1}\right]$ $=\mathbb{E}\left[\mathbbm{1}(V_{t}=1)\left(\bm{\theta}_{Q}^{\top}\bm{Q}_{t}^{(1)}-\bm{\theta}^{{}^{\prime}\top}_{Q}\bm{Q}_{t}^{(1)}\right)^{2}\mid\mathcal{H}_{t-1}\right]$
@@
Thus, combining the above display with Equation (H.2) gives the desired result in Equation (H.5). ∎
H.1.3 Proof of Lemma H.3

### v3

@@ I.2 Theoretical Guarantees and Regret Comparison
$\mathcal{K}^{(\texttt{non})}_{\bm{\theta},t}(f,f^{\prime})$ $=C_{0}\sum_{i=1}^{N}\mathbb{E}_{f}\left[f\left(\bm{O}_{i}^{(0)}\right)-f^{\prime}\left(\bm{O}_{i}^{(0)}\right)\right]^{2}$ (K.1)
$~~+C_{D}\sum_{\tau=1}^{t}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\left(f(\bm{O}_{\tau})-f^{\prime}(\bm{O}_{\tau})\right)^{2}\mathds{1}\left\{A_{\tau}=1,V_{\tau}=0\right\}\mid\mathcal{H}_{\tau-1}\right]$
@@
$\mathcal{K}^{(\texttt{lin})}_{f,t}\left(\bm{\theta},\bm{\theta}^{\prime}\right):=C_{D}\sum_{\tau=1}^{t}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathds{1}(V_{\tau}=1)\left(\bm{\theta}_{Q}^{\top}\bm{Q}_{\tau}^{(A_{\tau})}-\bm{\theta}^{{}^{\prime}\top}_{Q}\bm{Q}_{\tau}^{(A_{\tau})}\right)^{2}\mid\mathcal{H}_{\tau-1}\right].$ (K.2)
Then for any fixed policy $\pi$ and distribution $\mathbb{P}^{(t)}_{\bm{\theta},f,\pi}$ whose density is specified in Equation (J.16),
$\operatorname{KL}\left(\mathbb{P}^{(t)}_{\bm{\theta},f,\pi}\|\mathbb{P}^{(t)}_{\bm{\theta}^{\prime},f^{\prime},\pi}\right)\leq\mathcal{K}^{(\texttt{non})}_{\bm{\theta},t}(f,f^{\prime})+\mathcal{K}^{(\texttt{lin})}_{f,t}\left(\bm{\theta},\bm{\theta}^{\prime}\right).$ (K.3)
The proof of all the technical lemmas are deferred to Section K.1. Lemma K.2 controls $\mathcal{K}^{(\texttt{lin})}_{f,t}$, while $\mathcal{K}^{(\texttt{non})}_{\bm{\theta},t}(f,f^{\prime})$ is controlled by Equation (K.31) in the proof of Lemma K.5.
Lemma K.2.
@@
$\mathbb{P}\left(\bm{Q}=\mathbf{e}_{i}\right)=\frac{1}{d_{\texttt{lin}}}\quad\text{for }i\in[d_{\texttt{lin}}].$ (K.4)
@@
$\mathcal{K}^{(\texttt{lin})}_{f,t}\left(\bm{\theta},\bm{\theta}^{\prime}\right)=\frac{C_{D}t\left\|\bm{\theta}_{Q}-\bm{\theta}^{\prime}_{Q}\right\|_{2}^{2}}{2d_{\texttt{lin}}}.$ (K.5)
@@
For the distribution in Equation (J.16), the expected cumulative regret is given by
$\mathcal{R}_{T}(\bm{\theta},f)$ $=\sum_{t=1}^{T}\mathbb{E}^{(t-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[R^{\star}_{t}-R_{t}\mid\mathcal{H}_{t-1}\right]$ (K.6)
$=\sum_{t=1}^{T}\mathbb{E}^{(t-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathds{1}\left(V_{t}=1\right)\left(\bm{Q}^{\star}_{t}-\bm{Q}^{(A_{t})}_{t}\right)^{\top}\bm{\theta}_{Q}\mid\mathcal{H}_{t-1}\right]$
$~~+\frac{1}{2}\sum_{t=1}^{T}\mathbb{E}^{(t-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathds{1}\left(V_{t}=0\right)\left(f^{\star}(\bm{O}_{t})-f^{(A_{t})}(\bm{O}_{t})\right)\mid\mathcal{H}_{t-1}\right]$
@@
Lemma K.3 controls $\mathcal{R}_{T}^{(\texttt{lin})}(\bm{\theta},f)$.
Lemma K.3.
@@
$\sup_{\theta\in\Theta}\mathcal{R}^{(\texttt{lin})}_{T}(\bm{\theta},f)$ $\geq\frac{\sqrt{d_{\texttt{lin}}T}}{4}\exp\left\{-2C_{D}\right\}.$ (K.7)
@@
where $M>0$ is specified later in Equation (K.45). We index the bins by integers $k\in[M^{d_{\texttt{non}}}]$ via the mapping
@@
$\phi_{\beta}(\mathbf{o})=\begin{cases}(1-\|\mathbf{o}\|_{\infty})^{\beta}&\mbox{ if }0\leq\|\mathbf{o}\|_{\infty}\leq 1,\\ 0&\mbox{ if }\|\mathbf{o}\|_{\infty}>1.\end{cases}$ (K.8)
@@
$m=\lceil c_{m}M^{d_{\texttt{non}}}\rceil$ (K.9)
@@
$f_{\bm{\omega}}(\mathbf{o})=\sum_{j=1}^{m}\omega_{j}\varphi_{j}(\mathbf{o}),$ (K.10)
@@
$\varphi_{j}(\mathbf{o})=M^{-\beta}C_{\phi}\phi_{\beta}\left(2M[\mathbf{o}-\bm{b}_{j}]\right)\mathds{1}(\mathbf{o}\in B_{j})$ (K.11)
and $C_{\phi}>0$ is a constant specified in Equation (K.12). Note that for any $\mathbf{o}\in B_{j}$, the rescaled argument satisfies $2M(\mathbf{o}-\bm{b}_{j})\in\left[-1,1\right]^{d_{\texttt{non}}}$, so $\left\|2M(\mathbf{o}-\bm{b}_{j})\right\|_{\infty}\in[0,1]$, ensuring that $\varphi_{j}$ in Equation (K.11) is well-defined.
The function $f_{\bm{\omega}}$ is thus a linear combination of localized, smooth bump functions with disjoint supports. Lemma K.4 establishes that each $f_{\bm{\omega}}$ lies in $\mathcal{F}_{\beta,L}$ for a suitable choice of constant $L$.
Lemma K.4.
Suppose that $\beta\in(0,1]$. For any $\bm{\omega}\in\Omega_{m}=\{\pm 1\}^{m}$, the function $f_{\bm{\omega}}$ defined in Equation (K.10) belongs to the smoothness class $\mathcal{F}_{\beta,L}$ with $L=\beta 2^{\beta}C_{\phi}>0$.
@@
$C_{\phi}:=\frac{L}{\beta 2^{\beta}}.$ (K.12)
Based on this choice of packing set, Lemma K.5 controls $\mathcal{R}_{T}^{(\texttt{non})}(\bm{\theta},f)$.
Lemma K.5.
@@
$\sup_{f\in\mathcal{F}_{\beta,L}}\mathcal{R}^{(\texttt{non})}_{T}(\bm{\theta},f)=\Theta\left(TN^{-\frac{\beta}{2\beta+d_{\texttt{non}}}}\right),$ (K.13)
where $\mathcal{R}^{(\texttt{non})}$ is defined in Equation (K.6).
Taking Equations (K.13) with (K.19) into Equation (K.6), we have
@@
establishing the desired result in Equation (J.17).
K.1 Proof of Technical Lemmas
In this section, we present the proof of the Lemmas K.1-K.5 used in the proof of Theorem J.1. We will frequently use the Bretagnolle-Huber inequality given in the following theorem.
Theorem K.1 (Bretagnolle-Huber inequality).
@@
See Theorem 14.2 in Lattimore and Szepesvári (2020). ∎
K.1.1 Proof of Lemma K.1
@@
Recall the definition of $\mathbb{P}^{(t)}_{\bm{\theta},f,\pi}$ as stated in Equation (J.16). Eliminating the shared terms, it follows that
$\operatorname{KL}\left(\mathbb{P}^{(t)}_{\bm{\theta},f,\pi}\|\mathbb{P}^{(t)}_{\bm{\theta}^{\prime},f^{\prime},\pi}\right)$ $=\mathbb{E}^{(t)}_{\bm{\theta},f,\pi}\left[\log\frac{d\mathbb{P}^{(t)}_{\bm{\theta},f,\pi}}{d\mathbb{P}^{(t)}_{\bm{\theta}^{\prime},f^{\prime},\pi}}\right$] (K.14)
@@
For $\mathcal{K}_{1}$ in Equation (K.14), by the KL divergence assumption in Equation (J.13), we have
$\mathcal{K}_{1}=\sum_{i=1}^{N}\mathbb{E}_{f}\left[\log\frac{p^{(0)}_{f(\bm{O}^{(0)}_{i})}}{p^{(0)}_{f^{\prime}(\bm{O}^{(0)}_{i})}}\left(W^{(0)}_{i}\right)\right]\leq C_{0}\sum_{i=1}^{N}\mathbb{E}_{f}\left[f\left(\bm{O}^{(0)}_{i}\right)-f^{\prime}\left(\bm{O}^{(0)}_{i}\right)\right]^{2}.$ (K.15)
@@
$=\mathbb{E}^{(t)}_{\bm{\theta},f,\pi}\left[\log\frac{p_{\bm{\theta}^{\top}_{Q}\bm{Q}_{t}^{(A_{t})}}\left(R_{t}\right)}{p_{\bm{\theta}^{\prime\top}_{Q}\bm{Q}_{t}^{(A_{t})}}\left(R_{t}\right)}\mathds{1}\{V_{t}=0\}\right]+\mathbb{E}^{(t)}_{\bm{\theta},f,\pi}\left[\log\frac{p_{f^{(A_{t})}(\bm{O}_{t})}\left(R_{t}\right)}{p_{f^{\prime(A_{t})}(\bm{O}_{t})}\left(R_{t}\right)}\mathds{1}\{V_{t}=1\}\right]$
$=\sum_{a\in\mathcal{A}}\mathbb{E}^{(t-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathds{1}(A_{t}=a,V_{t}=0)\operatorname{KL}\left(\mathbb{P}_{\bm{\theta}_{Q}^{\top}\bm{Q}_{t}^{(a)}}\|\mathbb{P}_{\bm{\theta}^{{}^{\prime}\top}_{Q}\bm{Q}_{t}^{(a)}}\right)\mid\mathcal{H}_{t-1}\right]$
$~~+\sum_{a\in\mathcal{A}}\mathbb{E}^{(t-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathds{1}(A_{t}=a,V_{t}=1)\operatorname{KL}\left(\mathbb{P}_{f^{(a)}(\bm{O}_{t})}\|\mathbb{P}_{f^{\prime(a)}(\bm{O}_{t})}\right)\mid\mathcal{H}_{t-1}\right],$
where the last equality follows from the definition of KL divergence. Taking the above display and Equation (K.15) into Equation (K.14) yields that
$\operatorname{KL}\left(\mathbb{P}^{(t)}_{\bm{\theta},f,\pi},\mathbb{P}^{(t)}_{\bm{\theta}^{\prime},f^{\prime},\pi}\right)$ $=\mathcal{K}_{1}+\sum_{\tau=1}^{t}\sum_{a\in\mathcal{A}}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathds{1}(A_{\tau}=a,V_{\tau}=0)\operatorname{KL}\left(\mathbb{P}_{f^{(a)}(\bm{O}_{\tau})}\|\mathbb{P}_{f^{\prime(a)}(\bm{O}_{\tau})}\right)\mid\mathcal{H}_{\tau-1}\right]$ (K.16)
$+\sum_{\tau=1}^{t}\sum_{a\in\mathcal{A}}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathds{1}(A_{\tau}=a,V_{\tau}=1)\operatorname{KL}\left(\mathbb{P}_{\bm{\theta}_{Q}^{\top}\bm{Q}_{\tau}^{(a)}}\|\mathbb{P}_{\bm{\theta}^{{}^{\prime}\top}_{Q}\bm{Q}_{\tau}^{(a)}}\right)\mid\mathcal{H}_{\tau-1}\right]$
@@
$+C_{D}\sum_{\tau=1}^{t}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\left(f(\bm{O}_{\tau})-f^{\prime}(\bm{O}_{\tau})\right)^{2}\mathds{1}\left\{A_{\tau}=1,V_{\tau}=0\right\}\mid\mathcal{H}_{\tau-1}\right]$
$+C_{D}\sum_{\tau=1}^{t}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathds{1}(V_{\tau}=1)\left(\bm{\theta}_{Q}^{\top}\bm{Q}_{\tau}^{(A_{\tau})}-\bm{\theta}^{{}^{\prime}\top}_{Q}\bm{Q}_{\tau}^{(A_{\tau})}\right)^{2}\mid\mathcal{H}_{\tau-1}\right],$
where the last inequality follows from Assumption J.1 and Equation (J.15). Taking the definition of $\mathcal{K}^{(\texttt{non})}_{\bm{\theta},t}$ and $\mathcal{K}^{(\texttt{lin})}_{f,t}$ in Equations (K.1) and (K.2) into Equation (K.16) yields the desired bound in Equation (K.3). ∎
K.1.2 Proof of Lemma K.2
@@
By definition of $\mathbb{P}_{Q}$ in Equation (K.4),
@@
$\mathbb{E}\left[\mathds{1}(V_{t}=1)\left(\bm{\theta}_{Q}^{\top}\bm{Q}_{t}^{(A_{t})}-\bm{\theta}^{{}^{\prime}\top}_{Q}\bm{Q}_{t}^{(A_{t})}\right)^{2}\mid\mathcal{H}_{t-1}\right]$ $=\mathbb{E}\left[\mathds{1}(V_{t}=1)\left(\bm{\theta}_{Q}^{\top}\bm{Q}_{t}^{(1)}-\bm{\theta}^{{}^{\prime}\top}_{Q}\bm{Q}_{t}^{(1)}\right)^{2}\mid\mathcal{H}_{t-1}\right]$
@@
Thus, combining the above display with Equation (K.2) gives the desired result in Equation (K.5). ∎
K.1.3 Proof of Lemma K.3